20201002 SX SOUL Commissioning

00:50 David already have WUnit operating and OCAM chilled
  • He did use w_start_all
  • AUXLOOPS process did not start
  • Started it manually

01:39, 01:52 OSCO

[AOeng@soul-sxwfs OSCO]$ ls -l
total 16
-rw-rw-r--. 1 AOeng AOeng 48 Oct  2 01:40 20201002_013923.log
-rw-rw-r--. 1 AOeng AOeng 48 Oct  2 01:43 20201002_014248.log
-rw-rw-r--. 1 AOeng AOeng 48 Oct  2 01:53 20201002_015229.log
-rw-rw-r--. 1 AOeng AOeng 48 Oct  2 01:54 20201002_015418.log

01:39 The files above contain the AO Tracking numbers for the each of 3 sets taken during OSCO

[AOeng@soul-sxwfs OSCO]$ more 20201002_013923.log
20201002_013923
20201002_013936
20201002_013948

02:06 Third OSCO

-rw-rw-r--. 1 AOeng AOeng  3 Oct  2 02:06 20201002_020628.log
-rw-rw-r--. 1 AOeng AOeng 48 Oct  2 02:09 20201002_020851.log
  • 20201002_020628.log does not have any file names, thus the file size of 3 rather than 48. Not sure if data was taken but filenames lost, or there was some sort of error. I did not see the AdSec collect any data (no log lines). I'm guess the data were not collected.

02:30 Pal 10: el 74, provided Rmag 12.06, OCAM mag 13.4, bin 2, freq 914, 37 average pupil intensity, 5.4 photons/subapp
Gains pre Optical gain 12.54, 0.41, 1.05, gains post optical gain 2.10, 0.07 0.17 with Optical gain 0.18, NCPA trigger disabled

03:37 Slopes stopped being calculated after ResumeAO complete at 02:39:23
  • Jenny noticed the thelight was all concentrated in the upper right pupil
  • Barry put a pause in LUCI script
  • Check correction being applied on Mirror GUI and were all 0.0
  • GOPT was 0.17,
  • Gains were 2.10, 0.07, 0.17
  • clicked "show slopes" on OCAM gui
    • completely flat, 0.0s
    • no pupils shown
  • Opened loop on WFS Control GUI at 02:48:48
  • Slopes now showing up on OCAM GUI
    • 2 pupils are now seen, with 0.0 as background
  • Close loop at WFS Control GUI at 02:49:25
  • GOPT value is still correct (0.17)
  • Set gains to 2.10, 0.07, 0.17
  • turn on offloads to primary
  • all looks good
  • Barry resumed LUCI script
WfsArbitrator_130144|DEB|      1037|2020-10-02 02:39:15.768681|             MAIN > arbFunc offsetXYinternal done.
WfsArbitrator_130144|DEB|      1038|2020-10-02 02:39:23.619800|             MAIN > arbFunc doResumeLoop (, 'FLAOWFS', 'LUCIFER')
WfsArbitrator_130144|INF|      1039|2020-10-02 02:39:23.728468|             MAIN > Total counts/subap: 159.467226 - framerate: 915.0 - Flux: 1398
WfsArbitrator_130144|DEB|      1040|2020-10-02 02:39:23.730681|             MAIN > arbFunc doResumeLoop done.
WfsArbitrator_130144|DEB|      1041|2020-10-02 02:48:48.626188|             MAIN > arbFunc doStopLoop (, 'FLAOWFS', 'LUCIFER')
WfsArbitrator_130144|DEB|      1042|2020-10-02 02:48:48.630079|             MAIN > arbFunc doStopLoop done.
WfsArbitrator_130144|DEB|      1043|2020-10-02 02:49:23.956241|             MAIN > arbFunc doEnableDisturb (, 'FLAOWFS', 'LUCIFER', True, False)
WfsArbitrator_130144|DEB|      1044|2020-10-02 02:49:23.956965|             MAIN > arbFunc doEnableDisturb done.
WfsArbitrator_130144|DEB|      1045|2020-10-02 02:49:25.435404|             MAIN > arbFunc doCloseLoop (, 'FLAOWFS', 'LUCIFER')
WfsArbitrator_130144|DEB|      1046|2020-10-02 02:49:25.439305|             MAIN > arbFunc doCloseLoop done.
WfsArbitrator_130144|DEB|      1047|2020-10-02 02:53:05.407374|             MAIN > arbFunc doPauseLoop (, 'FLAOWFS', 'LUCIFER')
WfsArbitrator_130144|INF|      1048|2020-10-02 02:53:05.519531|             MAIN > Total counts/subap: 135.925153 - framerate: 915.0 - Flux: 1191
WfsArbitrator_130144|DEB|      1049|2020-10-02 02:53:05.577905|             MAIN > arbFunc doPauseLoop done.

JENNY's NOTES:

03:10 Rerotator in error, luci rotator at -3. We went to a negative rotator positions for the last 3 luci scripted images but just let it run. No tracking data taken...

3:16 Slewing to M52. Gains pre Optical gain 10.77, 0.71, 0.88, gains post optical gain 1.28 0.08 0.10 with Optical gain 0.124, NCPA trigger disabled
  • On first offset, the offset failed to resume. The baysides were almost in place. Doug moved the bayside and that kicked things in gear and reminded the baysides it was in place. Then the offset resumed.
  • Active guiding on sky offsets pulls the LUCIGS to the hotspot so when we return to source we are ~1-2" off and tend to safeskip. Suggested solutions: Split resume to start by raming tt for 1 sec then ho1 & ho2 once pulled on. OR once the loop is closed turn off active guiding until stopAO is run.
  • 3:39 NCPA turned on. LUCI images look great. Bin 2
4:28 GRB
  • the camera lens was pegged at 100, but pupils not able to center. Instrument rotator at 0 and increasing. Barry sending 90deg offset script.
  • The camera lens still pegged at luci rotator position of 90.
  • Luci rotator 270, elevation 65, camera lens X pos ~95, Y pos ~40 for popils centered in close loop, bin4, freq 491
  • AUto gain found tt2.7 ho12.7 but from plots tt2.5 and ho1~1 would be more reasonable so manually set. Aux loop process down and will not come back up. Process reset on the the other side, and GOPT would not apply. Optical gain of 1.000
  • No slopes again being calculated. Doug disabled offloads, moved the source back on to the cross hairs using the PSF GUI. Doug switched the binning to 3 and back to 4 to reset the BCU, and reapplied gains and reenabled offloafs and we are back to taking data.
5:31 GRB AOref star :
  • provided Rmag 16.61 (Actual catalog 15.06), OCAM mag 16.1, bin 4, freq 491, 26 average pupil intensity, 4.4 photons/subapp
  • Gains pre Optical gain 1.9, 1.0, 0.1,GOPT would not apply. Optical gain of 1.000, Aux loop process down and will not come back up
  • 20201002_053514 luci image luci1.20201009.0124,
  • 20201002_054350 luci image luci1.20201009.0128
  • 5:51 luci rotator at 99, camera lens X position skirting between +95-100 to keep pupils aligned. Y position ~ -10 (JUNK 20201002_0055238) TR#20201002_0055450
6:42 Safe skipping, right after a large offset. Paused on WFS gui but that did not pause on Adsec control gui, Pause AO on Adsec Control gui still saw safe skips. Open on WFS arb opened on Adsec. We are transiting at 86 deg. I think I caused that communication hiccup but restartin ghtre e wfs arb for good measure, then Barry resend preset
  • LUCI rotator at 156, camera lens X position is maxing out. Elevation is 88.5deg as we are closing. the estimated magnitude of the target this time is 17.2 when earlier was 16.2
  • LUCI rotator 244 deg pos y is -14 but X 100, bin4. Cannot center pupils
  • LUCI rotator 256 deg camera pos ~19, X is maxed at 100, pupi, diff 40, 0.02. Took coptical loop data during center pupils attempt 20201002_071954
  • Enrico and team on and witnessed
7:51 Performance testing Bin2
  • provided Rmag 12.8, OCAM mag 13.1, bin 2, freq 978, 47 average pupil intensity, 8.1 photons/subapp
  • Gains pre Optical gain Gains before optical gain 9.99, 0.52, 1.02, with optical gain applied 1.50, 0.08, 0.15
  • optical gain plots very flat and the lower end taken, ho1 flat out to 0.-5 to 2.0 but 0.52 taken. tt is flat 5-22 or so
NS test Eugina
  • Nothign intersting to report. provided Rmag 12 (Vmag 12.189, but we didn't look up any R mag references to be more precise), OCAM mag 13.4, bin 2, freq 904, 38 average pupil intensity, 6.5 photons/subapp. nice resolved image on LUCI
Performance star tests.
  • Star 11: El 61, Rot 188, sent Mag 15.3, OCAM16.6, bin 4, freq 435, average pupil intensity 29, 3.1 photons/subapp, closed vent doors and the estimated mag dropped to 17.1. Pointed North, away from the moon. With vent doors full open the OCAM estimated mag is 16.6, vent doors fully closed 17.1.... and changes as the vent doors are closing... New background needed?
  • 10:10 While sitting on star11, running MGM we walked off. Seeing had started to deteriorate from 1 to 1.3" but other than that there is no obvious reason. The loop remained closed but did not appear to be sending corrections. No slopes? We walked off so just resent preset.
  • Resent preset for star 11, vent doors closed this time: sent mag: 15.3, OCAM 16.7, DIMM 1.24", el 60, rot 209, bin 4 freq 457, 18 average pupil intensity, 3.0 photons/subapp. Seeing deteriorated to 1.5" on DIMM briefly then trending back to 1.1". Continuing Gopt tests followed by MGM tests.
  • the Optimize gain routine for bin 4 reoutinesly underestimates tt selecting 0.5 when clear minimum is 1.5-2.5, and overestimates ho1 setting to 2.7, when closer to 0.5 or 1.0. Its easier to just manually set.
  • 10:49 Agagin, loop is closed but no correction. I have disabled the offload and Enrico and team are taking TR data. The issue is on BCU2k, no slopes. Opening and reclosing returned the slopes.

  • Star 12: El 58, rotator 222, sent mag 15.3, OCAM mag 16.9, bin 4 freq 340, average pupil intensity 16, 1.6 photons/subapp. Gains pre GOPT 3.00, 1.00, 0.10, Gains after GOPT 0.34, 0.11, 0.01. Seeing 1.03 on the DIMM
  • just before finishing, the Gopt processes stopped on its own without dying.

  • 12:15 Star 13 el 53, rot 242,, sent mag 16.4, OCAM mag 17.3, bin 4, freq 266, average pupil intensity 15, 3.0 photons/subapp. Optimize gain overestimated ott and ho1 again. and luci1 out of sync after done and did not recognize startAO complete.
  • For this target, it was sent as bin3 but center pupils was taking too long with so few photons so once close enough we interrupted and just Check Flux to. kick into bin4 and centered in bin4. Perhaps luci did not recognize completed startAo because a step was incomplete in the sequence?
  • Seeing 0.7"

  • 12:27 AO248
  • Real time pupils disappeared during the close looop. No processes down but Cyled pupil check ocam2 process but real time pupils did not return. The team is investigating. Issue because too much flux from the background
-- DougMiller - 02 Oct 2020
Topic revision: r6 - 02 Oct 2020, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback